# Disentangled Interpretable Representation for Efficient Long-term Time Series Forecasting

### v1

@@ Abstract
Industry 5.0 introduces new challenges for Long-term Time Series Forecasting (LTSF), characterized by high-dimensional, high-resolution data and high-stakes application scenarios. Against this backdrop, developing efficient and interpretable models for LTSF becomes a key challenge. Existing deep learning and linear models often suffer from excessive parameter complexity and lack intuitive interpretability. To address these issues, we propose DiPE-Linear, a Disentangled interpretable Parameter-Efficient Linear network. DiPE-Linear incorporates three temporal components: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM). These components alternate between learning in the frequency and time domains to achieve disentangled interpretability. The decomposed model structure reduces parameter complexity from quadratic in fully connected networks (FCs) to linear and computational complexity from quadratic to log-linear. Additionally, a Low-Rank Weight Sharing policy enhances the model’s ability to handle multivariate series. Despite operating within a subspace of FCs with limited expressive capacity, DiPE-Linear demonstrates comparable or superior performance to both FCs and nonlinear models across multiple open-source and real-world LTSF datasets, validating the effectiveness of its sophisticatedly designed structure. The combination of efficiency, accuracy, and interpretability makes DiPE-Linear a strong candidate for advancing LTSF in both research and real-world applications. The source code is available at https://github.com/wintertee/DiPE-Linear.
Index Terms:
Time series analysis, Time series forecasting, Efficient neural network, Interpretable artificial intelligenceI Introduction
Long-term Time Series Forecasting (LTSF) is of significant importance in numerous domains, particularly in industrial applications such as cloud computing [1, 2], energy optimization [3, 4], and manufacturing [5, 6]. Recent advancements in deep learning, including Recurrent Neural Networks (RNNs) [7], Convolutional Neural Networks (CNNs) [8, 9, 10, 11], Transformers [12, 13, 14, 15, 16, 17, 18, 19, 20], Multilayer Perceptrons (MLPs)[21, 22, 23, 24, 25, 26] and Fully Connected linear models (FCs) [27, 28, 29] have demonstrated promising performance for LTSF tasks. With the advent of Industry 5.0 and the era of big data, time-series data processing faces three primary challenges: high dimensionality, high resolution, and regulatory industrial scenarios. As regulatory emphasis on carbon footprint and trustworthiness increases[30], developing an efficient and interpretable time-series forecasting method has emerged as a critical challenge.
Long-term forecasting for high-resolution time-series data remains challenging. Metrics like CPU load and network traffic in cloud environments are sampled at minute intervals, yet predictions must span days to support proactive resource allocation [31, 32, 20]. Similarly, advanced sensors in industrial settings generate vast minute-level data, requiring accurate long-term forecasting for equipment monitoring and operational optimization [33]. Existing methods face key limitations: RNNs struggle with parallelization [7], CNNs (e.g., TimesNet [11]) impose high computational costs, and Transformers, despite linear complexity improvements [16, 17], often overfit on long input sequences [27]. Fully connected models offer simplicity and competitive performance but suffer from quadratic complexity and parameter redundancy (Figs. 1(a), 1(b) and 1(c)). This raises the pivotal question: How can a model achieve parameter efficiency to effectively handle the challenges of long-term forecasting for high-resolution data?
In high-stakes industrial scenarios, interpretability are indispensable requirements for achieving trustworthy AI [34, 35]. In this context, domain experts must understand the model to trust its predictions. While leading deep learning models offer impressive precision, their black-box nature significantly undermines user trust. Fully connected linear models, by contrast, provide better interpretability by design. However, it remains challenging to fully and directly understand the features encoded by these models through their two-dimensional weight matrices, whether represented in the time domain (Figs. 1(a) and 1(b)) or frequency domain (Fig. 1(c)). This introduces the second key question: How can a model ensure interpretability, enabling domain experts to understand and trust its predictions in high-stakes industrial applications?
To address these challenges, we revisit the performance of FCs in LTSF, seeking inductive biases to reduce complexity and enhance interpretability. Figs. 1(a) and 1(b) reveals that highly correlated weights along the diagonal suggest parameter redundancy in the time domain, while the high noise in off-diagonal regions in the frequency domain as shown in Fig. 1(c) indicates the independence of frequency components. This motivates our frequency independence hypothesis. Furthermore, time series often exhibit varying contributions of input time points to future predictions; for example, the most recent data typically plays a greater role in predicting future values. Thus, modeling the temporal significance of input sequences is crucial. Similarly, different frequency components contribute unequally[36], with significant frequencies often possessing higher signal-to-noise ratios and deserving more model attention. Finally, in multivariate time series, variable dependencies must be captured in an explainable manner, as variables may either share patterns or act independently.
Based on the aforementioned motivations, we propose a novel Disentangled interpretable and Parameter-Efficient Linear model (DiPE-Linear) for long-term time series forecasting. As a deep linear model, it is structured into three interpretable linear sub-modules: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM). The SFA module extracts and enhances prominent frequencies in the frequency domain, and the STA module selects input time points in the time domain that are strongly correlated with the predicted sequence. Meanwhile, the IFM module assumes independence across different frequencies and directly maps the input series to the output series at each frequency component in frequency domain. By leveraging disentangled representations in both the time and frequency domains, the model achieves reduced complexity with linear parameter growth and log-linear computational scaling. Additionally, as shown in Fig. 1(d), these representations enhance domain experts’ trust by providing clear insights into the learned features and the model’s reasoning process.
These sub-modules employ a low-rank weight-sharing policy to efficiently model inter-variate relationships, reducing model complexity while maintaining strong performance. We further introduce a novel loss function, SFALoss, which is defined as a sum of time-domain loss and frequency-domain loss weighted by frequency-domain attention in the SFA module, guiding the model to prioritize significant frequencies with less noise.
Additionally, by avoiding non-linearity within these modules, our model remains a linear model while achieving comparable or SOTA performance in LSTF with significantly fewer parameters, as shown in Fig. 2. This lightweight design enhances model’s interpretability and establishes a more reliable and efficient prediction pipeline, offering valuable insights for future research.
@@
We propose DiPE-Linear, a deep linear model that achieves efficiency, significantly reducing parameter complexity from quadratic to linear and computational complexity from quadratic to log-linear.
Disentangled representation of model components in both time and frequency domain offers superior interpretability, thereby enhancing its trustworthiness for industrial applications.
Extensive experiments demonstrate that our method outperforms existing approaches, offering superior performance, efficiency, and interpretability.
II Related Work
II-A Efficient Architectures for Long-Term Forecasting
Long-term time series forecasting (LTSF) requires models capable of efficiently handling long sequences. Early Transformer-based models [12, 13, 14, 15, 16, 17] progressively reduced computational complexity from to . However, due to the inherently complex structure of Transformers, these models still contain a large number of parameters and are highly prone to overfitting when dealing with excessively long input sequences in LTSF tasks[27].
@@ II-A Efficient Architectures for Long-Term Forecasting
DLinear[27] was the first to propose using a simple single-layer fully connected network to replace the complex Transformer structure. Although the fully connected layer has a computational complexity of , its simple structure results in significantly lower complexity compared to other Transformer-based models. Nevertheless, DLinear still suffers from notable parameter redundancy. FITS[29] addressed this by employing low-pass filters to discard high-frequency information in the frequency domain, drastically reducing the number of parameters in the fully connected network to the scale of 10k. SparseTSF[37], on the other hand, introduced Cross-Period Sparse Forecasting, which achieves a parameter count on the scale of 1k through downsampling and cross-period parameter sharing. However, both FITS and SparseTSF sacrifice accuracy to achieve parameter reduction[38, 37]. Further exploration into addressing parameter redundancy in fully connected models remains an open research question.
II-B Exploiting Frequency-Domain Characteristics
Recent advancements in time series forecasting have increasingly leveraged frequency-domain knowledge to improve performance. One prominent approach incorporates frequency-domain information as auxiliary features to enhance the modeling of long-term dependencies. For example, FEDformer[17] introduces a DFT-based frequency-enhanced attention mechanism, where attention weights are computed from the spectra of queries and keys, with the weighted sum calculated directly in the frequency domain. Similarly, FiLM[39] applies Fourier analysis to retain historical information while filtering out noise. Alternatively, some methods operate entirely in the frequency domain. FreTS[40] employs a frequency-domain MLP to capture both channel-wise and temporal dependencies, while FITS[29] adopts a fully connected network specifically designed for spectral representations.
Another promising direction incorporates frequency-domain loss functions into time-domain training objectives to capture spectral patterns more effectively. For example, FTMixer [26] and FreDF [41] add frequency-domain loss components on top of time-domain losses, enhancing the model’s ability to align with the ground truth in the frequency domain.
II-C Channel-Independence for Multivariate Time Series
DLinear[27] was the first to introduce the Channel-Independent strategy for multivariate time series forecasting (the terms ”channel” and ”variate” are used interchangeably throughout this paper for simplicity). Although this approach ignores causal relationships between variables, it achieves significant performance improvements. PatchTST[19] further validated this idea in Transformer-based models, demonstrating that independently processing variables while sharing the same model weights across variables can be highly effective.
However, RLinear[28] shows that for variables with distinct periodic characteristics, using independent weights for each variable can further enhance performance. Subsequent research has focused on finding an optimal balance between weight sharing and weight independence. For instance, MLinear[42] dynamically adjusts between Weight Sharing and Weight Independent strategies based on the temporal semantics of different sequences, while CSC[43] reassigns weight-variable mappings during the validation phase to achieve variable clustering. Nevertheless, these methods come with limitations. MLinear requires calculating both Weight Sharing and Weight Independent outputs before combining them, which increases computational overhead. CSC, on the other hand, relies on a strong prior assumption that variables clustered together should share the same weights, which may not hold universally.
III Method
@@
DiPE-Linear is constructed from several essential components. We begin by outlining the preliminaries related to TSF and the Fourier Transform, providing the foundation for our approach. Next, we present the architecture of DiPE-Linear, detailing its three core modules: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM), under assumption that the input series is univariate. Then, we describe how the Low-rank Weight Sharing method generalizes those modules to multivariate time series. Lastly, we introduce SFALoss, a novel loss function that combines weighted losses in both the frequency and time domains, guiding the model to prioritize informative patterns while enhancing predictive performance.
III-A Preliminaries
III-A1 Problem Definition
@@
The multivariate TSF problem can be formally defined as follows: Given a historical time series , where is the number of variates and is the look-back length of input time series, the objective is to predict the future time series , where is the forecasting horizon.
III-A2 Discrete Fourier Transform
Time series data can be viewed as discrete samples of a continuous signal over time. In many signal processing and time-series forecasting applications, the Discrete Fourier Transform (DFT) is a commonly used method to analyze the frequency components of such discrete signals. For a discrete-time sequence of length , the DFT transforms the time-domain signal into its frequency-domain representation. This transformation is particularly useful for identifying underlying periodic patterns or extracting frequency components that contribute to the signal.
The DFT of a time series is defined as:
Once the signal is represented in the frequency domain, various operations, such as filtering or modifying specific frequency components, can be applied. The inverse Discrete Fourier Transform (iDFT) allows for transforming the modified frequency-domain signal back to the time domain. The iDFT is expressed as:
This transformation provides the flexibility to manipulate signals in the frequency domain and revert them back to the time domain for further analysis or application. The ability to move between domains seamlessly is essential for efficient signal processing.
III-A3 Duality Between Time and Frequency Domains
The DFT exhibits a fundamental property known as duality between the time and frequency domains. This duality manifests as convolution in the time domain being equivalent to multiplication in the frequency domain, and vice versa. For two time-domain signals, and , their convolution:
has a frequency-domain equivalent:
Pointwise multiplication in the time domain, on the other hand, corresponds to convolution in the frequency domain. This relationship simplifies many operations, particularly in tasks where convolution is computationally intensive in the time domain.
Thus, the DFT provides a powerful tool for analyzing time series in the frequency domain, leveraging the duality between time and frequency domains to simplify complex operations such as convolution. The ability to transition between domains efficiently enables advanced signal processing techniques for both single-channel and multichannel data.
III-B Static Frequential Attention
Extracting frequencies is a common practice, as previous FITS [29] employed a low-pass filter to eliminate high-frequency signals, while FedFormer [17] randomly selects a subset of frequencies as key features. However, such strong prior assumptions may lead the model to discard useful information. Therefore, we aim for the model to learn which frequency components to focus on.
@@ III-B Static Frequential Attention
Specifically, to ensure the model’s capability to select and amplify significant frequencies, we introduce SFA module as a frequency-domain filter. Additionally, to preserve the temporal structure of the series and avoid interference with subsequent temporal feature extraction, we constrained this filter to be a zero-phase filter. The SFA operates by first applying real Fast Fourier Transform (rFFT) to transform the time-domain data into the frequency domain. In this domain, a learned filter is applied via element-wise multiplication, selectively amplifying or suppressing specific frequency components. As a zero-phase filter, it exclusively modifies the amplitude of the signal while preserving the original phase information. Finally, inverse rFFT (irFFT) is used to reverse the filtered data back to the time domain. Specifically, the enhanced signal is given by
@@
III-C Static Temporal Attention
Similarly, to enable the model to have the capability of capturing important input time points in the temporal domain, we introduce the STA as a second component of our model. This module enables the input time series to undergo element-wise multiplication with a learned temporal attention map, effectively assigning appropriate importance to relevant historical time points and thereby enhancing the model’s ability to capture temporal dependencies. The processed signal is thus given by
@@ III-C Static Temporal Attention
III-D Independent Frequential Mapping
While both the SFA and STA modules function as part of the input series pre-processing, the IFM module directly maps the input historical series to the forecasted output. To process the time series as a whole, leveraging the spectrum is a natural approach, as it encapsulates all the information about the entire sequence and enables better handling of periodic components. Properly and efficiently processing the spectrum remains a key challenge. Fig. 1(c) illustrates that the two-dimensional matrix used by FITS [29] to handle the spectrum is sparse, with only the diagonal weights being significant. This suggests that, for a time series, there is relatively strong independence between different frequencies, allowing each frequency to be handled separately.
We thus treat each frequency independently for two key reasons: (1) Apart from harmonic frequencies, there is generally little correlation observed between different frequencies in time series data.[41] (2) Although harmonic frequencies tend to exhibit strong correlations, we avoid explicitly modeling these inter-frequency dependencies to mitigate potential multicollinearity in linear regression.
Therefore, for each frequency component, we apply a complex-valued multiply-accumulate operation independently. To transform the input space to the output space , we apply zero-padding to the input series, extending its length to before performing the rFFT. At the end, the last values are extracted and used as the predicted output. The forecasted output is thus defined as:
where is the complex-valued weight and is the complex-valued bias.
This process can be equivalently described as a 1D convolution in time domain, as explained in equation 3 and 4:

### v2

@@ Abstract
Long-term Time Series Forecasting (LTSF) is crucial across various domains, but complex deep models like Transformers are often prone to overfitting on extended sequences. Linear Fully Connected (FC) models have emerged as a powerful alternative, achieving competitive results with fewer parameters. However, their reliance on a single, monolithic weight matrix leads to quadratic parameter redundancy and an entanglement of temporal and frequential properties. To address this, we propose DiPE-Linear, a novel model that disentangles this monolithic mapping into a sequence of specialized, parameter-efficient modules. DiPE-Linear features three core components: Static Frequential Attention to prioritize critical frequencies, Static Time Attention to focus on key time steps, and Independent Frequential Mapping to independently process frequency components. A Low-rank Weight Sharing policy further enhances efficiency for multivariate data. This disentangled architecture collectively reduces parameter complexity from quadratic to linear and computational complexity to log-linear. Experiments on real-world datasets show that DiPE-Linear delivers state-of-the-art performance with significantly fewer parameters, establishing a new and highly efficient baseline for LTSF. Our code is available at https://github.com/wintertee/DiPE-Linear/
1 Introduction
Long-term Time Series Forecasting (LTSF) plays a crucial role in a wide array of fields, such as weather forecasting [1] and industrial manufacturing [27], making it a prominent focus of research in real-world applications. However, LTSF presents significant challenges due to the need for processing extended look-back window and forecasting horizon, which increases both the complexity of the forecasting task and the dimensionality of the model. Recent advancements in deep learning, such as Convolutional Neural Networks [12], and Transformers [20, 14, 11], have shown promise in tackling these challenges. However, as sequence length increases, the number of model parameters grows rapidly, significantly heightening the risk of overfitting [23].
| (a) | DLinear [23] |
| MSE=0.399 | |
| #Params=18K |
Trend
Seasonal
IR
| (b) | FITS[21] |
| MSE=0.385 | |
| #Params=6.4K |
Amplitude
Phase
IR
| (c) | DiPE-Linear (Ours) |
| MSE=0.384 | |
| #Params=0.7K |
SFA
STA
IFM
IR
In response to this overfitting challenge, Linear Fully Connected (FC) models have gained renewed attention. Recent research has shown that such FC models can match or even outperform complex nonlinear models, while requiring substantially fewer parameters [23, 7, 21, 10]. However, despite addressing the parameter excess of nonlinear models, these FC models are often monolithic. They attempt to learn complex temporal dynamics using a single, dense weight matrix . This all-in-one design, which simultaneously captures frequential properties, temporal dependencies, and the input-output mapping, introduces its own critical issues: it not only leads to significant parameter redundancy, evidenced by its quadratic complexity, but also obscures the model’s inner working mechanisms.
The inefficiency of this monolithic design can be observed directly. For instance, the weight matrix of a trained DLinear[23] model (Fig. 1a) exhibits distinct horizontal and diagonal stripes. These stripes represent the same pattern repeating at different periodic lags, revealing a massive parameter redundancy—the model is inefficiently storing the same temporal dependency multiple times. Meanwhile, the FITS[21] off-diagonal amplitude and phase visualization (Fig. 1b) appears noisy. These visualizations suggest that while FC models attempt to learn structural properties, they do so implicitly and inefficiently. We identify this core problem as entanglement: the properties are wastefully embedded within a single, large, dense matrix.
This motivates our core idea: to replace the monolithic matrix with an explicit factorization, or disentanglement, into a sequence of specialized, interpretable, and parameter-efficient linear modules. We disentangle the core temporal dynamics into three key priors: (1) Frequential Filtering: In recognition of the fact that time series data are typically decomposed into trend and seasonal components [20], we incorporate the Static Frequential Attention (SFA) module to prioritize predictable frequencies. (2) Temporal Importance: Not all historical time steps are equally important. This motivates the Static Time Attention (STA) module to focus on pivotal time steps. (3) Independent Mapping: To model these periodicities holistically and efficiently through a time-invariant mapping (thus avoiding the explicit redundancy seen in Fig. 1), we propose the Independent Frequential Mapping (IFM) approach. Furthermore, to efficiently capture inter-variable dependencies [15] in the multivariate setting, we introduce a (4) Low-rank Weight Sharing policy. Our explicit factorization (Fig. 1c) achieves a cleaner, more efficient representation, directly addressing the redundancy observed in Fig. 1a and 1b.
Building upon the above-mentioned motivation, we propose a novel Disentangled Parameter-Efficient Linear model (DiPE-Linear) for LTSF. In this work, "disentangled" refers to the model’s architecture which factorizes the complex, monolithic FC weight matrix into a sequence of specialized, interpretable, and parameter-efficient linear modules (SFA, STA, and IFM). Each module addresses a distinct aspect of temporal modeling (frequential filtering, temporal importance weighting, and frequential mapping, respectively). These linear sub-modules collectively reduce the FC’s quadratic parameter complexity to linear complexity and quadratic computational complexity to log-linear. This lightweight, fully linear design achieves state-of-the-art performance with minimal parameters, as shown in Fig. 1 and 2, establishing an efficient and reliable pipeline for time series forecasting.
@@ 1 Introduction
We propose DiPE-Linear, a novel disentangled parameter-efficient linear model for LTSF. Its innovative design, motivated by an analysis of entanglement in existing FCs, features specialized sub-modules that improve performance and offer clear interpretability.
The proposed DiPE-Linear significantly reduces the parameter complexity (from quadratic to linear) as well as computational complexity (to log-linear) without sacrificing accuracy.
Extensive experiments demonstrate that DiPE-Linear achieves performance comparable to or surpassing other state-of-the-art models, validating its effectiveness and efficiency.
2 Related Work
2.0.1 Efficient Long-Term Forecasting.
@@ 2 Related Work
LTSF demands models that handle long sequences efficiently. Early Transformer-based models [5, 4, 24, 20, 26] reduced computational complexity from to , but their complex structure leads to a large number of parameters, causing overfitting on long inputs [23]. DLinear [23] proposed a simpler, single-layer fully connected network with complexity, offering lower overall complexity than Transformer-based models, though it still suffers from parameter redundancy. FITS [21] reduced parameter count by applying low-pass filters in the frequency domain, while SparseTSF [10] introduced Cross-Period Sparse Forecasting to achieve further parameter reduction. However, both methods trade off accuracy to reduce parameters [16, 10], and addressing parameter redundancy in FC models remains an open challenge.
2.0.2 Exploiting Frequency-Domain Characteristics.
Recent advances in time series forecasting have leveraged frequency-domain techniques to enhance performance. FEDformer [26] introduced a DFT-based frequency-enhanced attention mechanism, while FiLM [25] used Fourier analysis to retain historical information and reduce noise. Other methods, such as FreTS [22] and FITS [21], work entirely in the frequency domain, with FreTS using a frequency-domain MLP for both temporal and channel-wise dependencies. Additionally, frequency-domain loss functions have been integrated into time-domain objectives. For example, FTMixer [8] and FreDF [17] use frequency-domain loss components to improve alignment with the ground truth.
2.0.3 Channel Dependence for Multivariate Forecasting.
DLinear [23] introduced the Channel-Independent strategy for multivariate time series forecasting, which, despite ignoring causal relationships, led to notable performance gains. PatchTST [14] further validated this approach in Transformer-based models, showing that independently processing variables while sharing model weights across them can be effective. RLinear [7] demonstrated that using independent weights for variables with distinct periodicities can further improve performance. Subsequent research has sought an optimal balance between weight sharing and independence. MLinear [6] dynamically adjusts strategies based on temporal semantics, while CSC [15] reassigns weight-variable mappings during validation to cluster variables. However, MLinear introduces extra computational overhead, and CSC relies on assumptions about weight sharing that may not always hold.
3 Method
@@
We first define the multivariate LTSF problem, then describe the DiPE-Linear architecture, focusing on its three core modules: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM), assuming univariate time series. We then explain how the Low-rank Weight Sharing method extends these modules to multivariate time series. Finally, we introduce SFALoss, a loss function that combines weighted losses in both frequency and time domains to prioritize predictable frequencies and improve model performance.
3.1 Problem Definition
@@
The multivariate LTSF problem is formally defined as follows: Given a historical time series , where denotes the number of variates and represents the look-back length, the goal is to predict the future time series , where is the forecasting horizon.
3.2 Static Frequential Attention
@@
We introduce the SFA module as a frequency-domain filter to enable the model to selectively enhance and amplify relevant frequencies. Additionally, to preserve the temporal structure of the series and avoid interference with subsequent temporal feature extraction, we constrained this filter to be a zero-phase filter. The SFA operates by first applying real Fast Fourier Transform (rFFT) to transform the time-domain data into the frequency domain. In this domain, a learned filter is applied via element-wise multiplication, selectively amplifying or suppressing specific frequency components. As a zero-phase filter, it exclusively modifies the amplitude of the signal while preserving the original phase information. Finally, inverse rFFT (irFFT) is used to reverse the filtered data back to the time domain. Specifically, the enhanced signal is given by
@@
3.3 Static Temporal Attention
We apply SFA before STA based on the hypothesis that the model should first isolate the most predictable, core frequencies of the signal (SFA), and then identify the critical time steps within that filtered signal (STA). Reversing the order might cause the STA module to focus on temporally salient but high-frequency noise.
Following this, to enable the model to have the capability of capturing important input time points in the temporal domain, we introduce the STA as a second component of our model. This module enables the input time series to undergo element-wise multiplication with a learned temporal attention map, effectively assigning appropriate importance to relevant historical time points and thereby enhancing the model’s ability to capture temporal dependencies. The processed signal is thus given by
@@ 3.3 Static Temporal Attention
3.4 Independent Frequential Mapping
While the SFA and STA modules are part of the input series pre-processing, the IFM module directly maps the input historical series to the forecasted output under the assumption of independence between different frequencies.
We acknowledge that this assumption is a theoretical simplification, as real-world signals often exhibit complex inter-frequency dependencies (e.g., harmonics [17]). However, we adopt this assumption as a deliberate design choice. Explicitly modeling these intricate dependencies within a linear framework could increase the risk of multicollinearity and overfitting.
Therefore, our approach acts as an effective inductive bias. By enforcing frequency-wise independence, the IFM module learns a 1D convolution (Eq. 6) that is inherently time-invariant and phase-preserving (relative to each frequency component), which are desirable properties for modeling stationary seasonal patterns. This compels the model to focus on the primary predictive information within each frequency component rather than capturing potentially spurious cross-frequency correlations. As our extensive experiments on diverse datasets (see Section 4) demonstrate, this design choice consistently leads to state-of-the-art performance, suggesting it enhances model generalization and robustness rather than limiting its accuracy.
Following this assumption, for each frequency component, we apply a complex-valued multiply-accumulate operation independently. To transform the input space to the output space , we apply zero-padding to the input series, extending its length to before performing the rFFT. At the end, the last values are extracted and used as the predicted output. The forecasted output is thus defined as:
where is the complex-valued weight of the IFM module, and is the complex-valued bias.
According to the Convolution Theorem, this process can be equivalently represented as a 1D convolution in the time domain: